Solve the system by substitution.
x + y = -7
x - y = 13

Answer by fractalier(6550)

Okay, let us solve the first one for x and substitute that value into the second one...like this...
x + y = -7
x = -y - 7
and
-y - 7 - y = 13
-2y - 7 = 13
-2y = 20
y = -10
and
then x = 3